Determine the distance y to the centroid of the beams cross-sectional area; moment of inertia about the x axis then find the 6 in 2 in. 4 in. 1 in. 1 in.
Answer #1

al = 6x2 = 12 (n° sin Ain A= 201·N tin 12 + 4 y= 2.2 in グース